OpenStudy (anonymous):

A square piece of fabric has an area of 595 cm2. To the nearest tenth, what is the length of 1 side of the fabric? A. 21.3 cm B. 23.2 cm C. 24.4 cm D. 25.6 cm

OpenStudy (anonymous):

@JackofallTradez

OpenStudy (campbell_st):

take the square root of 595... that will give the side length remember the area of a square is \[A = l^2 ~~~so~~~~\sqrt{A} = l\]

OpenStudy (jackofalltradez):

Since it's a square, both sides are identical. This can be represented by x (side x) * x (side x) = 592 cm^2. Now, this simplifies to x^2 = 592 cm^2. Now, solve for x

OpenStudy (jackofalltradez):